Two identical conducting spherese are placed with their centers 45 cm apart one is given a charge of +12 x 10^-9 C and the other is given a charge of -18 x 10^-9 C. the spheres are connected by a conducting wire. After equilibrium has occurred, find the electric force between the two spheres



1
Expert's answer
2022-02-22T04:41:22-0500

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}\\
=9*10^9*\frac{12* 10^{-9}*18* 10^{-9}}{0.45^2}=9.6*10^{-6}\:\rm N"